A 19.0 m -long copper wire, 2.80 mm in diameter including insulation, is tightly wrapped in a single layer with adjacent coils touching, to form a solenoid of diameter 2.70 cm (outer edge).

What is the length of the solenoid? (M)

What is the field at the center when the current in the wire is 20.0 A ? (M)
Answer #1

for the length of the solenoid

number of turns = 19/(2pi * (0.027/2))

number of turns = 224

length of solenoid = 224 * 2.8 mm

length of solenoid = 627.2 mm = 0.627 m

the length of solenoid is 0.627 m

for the field at the center

field at the center = u0 * I * N/L

field at the center = 4pi *10^-7 * 224 * 20/0.627

field at the center = 0.00897 T

the magnetic field at the center is 8.97 mT